OpenStudy (alinaa1024):

A nick on the edge of a CD rotates to (−6, 5) during one song when represented graphically. What is the sine value of this function?

OpenStudy (thereisnousername):

I would first try to graph the point :  sin Angle = opposite / hypotenuse the opposite side has length 5 we need the length of the hypotenuse. Use pythagoras to figure out the length of the hypotenuse. i.e. a^2 + b^2 = c^2 can you find the length of the hypotenuse ?

OpenStudy (alinaa1024):

I don't understand what would I have to do to find the hypotenuse?

OpenStudy (alinaa1024):

@ThereisnoUsername

OpenStudy (thereisnousername):

 A and B are practically X and Y, C is unknown yet, so plug in the X and Y values, what do you get?

OpenStudy (alinaa1024):

can you set it up with the variables having numbers?

OpenStudy (thereisnousername):

What is the X value,

OpenStudy (alinaa1024):

-6

OpenStudy (thereisnousername):

okay, and the Y?

OpenStudy (alinaa1024):

5

OpenStudy (thereisnousername):

okay, so now plug that in for A and B : (5)^2 + (-6)^2 = c^2, and what do you get?

OpenStudy (thereisnousername):

you could also look at it vice versa: (-6)^2+(5)^2=c^2

OpenStudy (alinaa1024):

61?

OpenStudy (thereisnousername):

yes, okay so now we have: 61=c^2 , and simplify: 7.81 = c

OpenStudy (alinaa1024):

is that equivalent to -sqrt of 61?

OpenStudy (thereisnousername):

no the square root of 61 is 61, 5/7.81 = Sin Angle so divide them and what do you get?

OpenStudy (alinaa1024):

0.640

OpenStudy (thereisnousername):

okay so I think the answer is: 5 times square root 61 over 61

OpenStudy (alinaa1024):

thank you! do you mind helping with more ?

OpenStudy (thereisnousername):

sure open a new question and tag me,

OpenStudy (alinaa1024):

okie dokie